# Generalized Brillouin Zone Fragmentation

### v1

@@ Abstract
Importantly, under usual circumstances with conventional GBZ, $S$ changes negligibly with the system size, shifting by less than 1% in the Hatano-Nelson model shown. This is because $S$ is insensitive to the discretization of the eigenvalue points, as long as the OBC spectrum does not change qualitatively with system size. To have a sense of what can change $S$, we observe that $S$ approximately doubles when the non-Hermitian asymmetry $\kappa$ is changed from $0.2$ to $2$, such that the inner OBC spectrum is much further away from the PBC spectral ring. As such, by measuring $S$, one can obtain a qualitative sense of how much the PBC and OBC spectra differ, independently of its implementation details i.e. resolution or lattice discretization.
Figure S16: PBC and OBC spectral distributions of the Hatano-Nelson model $e^{\kappa}z+e^{-\kappa}/z$ (a-d) and a random instance of our $b=3$-component random hopping model $H_{\text{rand}}$ [Eq. S40], computed from numerical data via Eq. S51 with Gaussian dot size $\delta=1/3$ and resolution $r=100$. These parameters should not be confused with the system size $N$, which controls the number of complex eigenvalues. In (a-d), $S$ is shown to depend only on $\kappa$ but not NN, since the relative entropy compares the overall shapes of the spectra, rather than their discretization details. The cleanly distinct OBC and PBC spectra in (a-d) contrasts with those of (e-f), which are characteristic of a fragmented GBZ.
@@
S5.2.2 Variation of spectra with system length $N$ and transverse spacing $g$
Finally, in Figs. S19 and S20, we vary the number of unit cells $N$ per row, such as to check whether our photonic crystal exhibits the spectral scaling characteristic of GBZ fragmentation. Also, in order to control the influence of inter-chain coupling, we vary the transverse spacing between the dielectric rows through $g$ [Fig. S17(b)]. All other parameters will be kept fixed relative to lattice constant $a$, which does not affect the physics due to the scale invariance of Maxwell’s equations.
Most saliently, for small spacing $g<1$, the PBC and OBC spectra differ much more in the aligned [Fig. S19] than in the anti-aligned [Fig. S20] structure. This is particularly prominent at larger number of unit cells $N$, i.e. for $g=0.4$, $N=100$. Ostensibly, this is because the antagonistic NHSE mechanism is strongest when the chains have a long runway, and when the NHSE channels are oppositely directed.
Note that due to the complex tangle of effective hoppings that must exist in the effective model, some antagonistic NHSE still exists even in the aligned cases, such that the PBC and OBC spectra generally becomes more similar at larger $N$. Another subtlety is that as the spacing $g$ gets larger, the PBC and OBC spectra eventually becomes more similar. This is because while increasing $g$ weakens the NHSE cancellation initially, large $g$ also leads to more compact spectral dispersion (flatter bands) in the complex plane, which are less affected by the boundaries.

### v2

@@ Abstract
Importantly, under usual circumstances with conventional GBZ, $S$ changes negligibly with the system size, shifting by less than 1% in the Hatano-Nelson model shown. This is because $S$ is insensitive to the discretization of the eigenvalue points, as long as the OBC spectrum does not change qualitatively with system size. To have a sense of what can change $S$, we observe that $S$ approximately doubles when the non-Hermitian asymmetry $\kappa$ is changed from $0.2$ to 22, such that the inner OBC spectrum is much further away from the PBC spectral ring. As such, by measuring $S$, one can obtain a qualitative sense of how much the PBC and OBC spectra differ, independently of its implementation details i.e. resolution or lattice discretization.
Figure S16: PBC and OBC spectral distributions of the Hatano-Nelson model $e^{\kappa}z+e^{-\kappa}/z$ (a-d) and a random instance of our $b=3$-component random hopping model $H_{\text{rand}}$ [Eq. S40], computed from numerical data via Eq. S51 with Gaussian dot size $\delta=1/3$ and resolution $r=100$. These parameters should not be confused with the system size NN, which controls the number of complex eigenvalues. In (a-d), $S$ is shown to depend only on $\kappa$ but not NN, since the relative entropy compares the overall shapes of the spectra, rather than their discretization details. The cleanly distinct OBC and PBC spectra in (a-d) contrasts with those of (e-f), which are characteristic of a fragmented GBZ.
@@
S5.2.2 Variation of spectra with system length NN and transverse spacing $g$
Finally, in Figs. S19 and S20, we vary the number of unit cells NN per row, such as to check whether our photonic crystal exhibits the spectral scaling characteristic of GBZ fragmentation. Also, in order to control the influence of inter-chain coupling, we vary the transverse spacing between the dielectric rows through $g$ [Fig. S17(b)]. All other parameters will be kept fixed relative to lattice constant $a$, which does not affect the physics due to the scale invariance of Maxwell’s equations.
Most saliently, for small spacing $g<1$, the PBC and OBC spectra differ much more in the aligned [Fig. S19] than in the anti-aligned [Fig. S20] structure. This is particularly prominent at larger number of unit cells NN, i.e. for $g=0.4$, $N=100$. Ostensibly, this is because the antagonistic NHSE mechanism is strongest when the chains have a long runway, and when the NHSE channels are oppositely directed.
Note that due to the complex tangle of effective hoppings that must exist in the effective model, some antagonistic NHSE still exists even in the aligned cases, such that the PBC and OBC spectra generally becomes more similar at larger NN. Another subtlety is that as the spacing $g$ gets larger, the PBC and OBC spectra eventually becomes more similar. This is because while increasing $g$ weakens the NHSE cancellation initially, large $g$ also leads to more compact spectral dispersion (flatter bands) in the complex plane, which are less affected by the boundaries.